2(5-4x)+x=3(3x-11)
Found 2 solutions by wuwei96815, checkley71:
Answer by wuwei96815(245)

2(5-4x) + x = 3(3x-11)

First, remove the parentheses.
10 -8x +x = 9x -33

Next, combine like terms.
-8x + x - 9x = -33 -10
-16x = -43

Finally, solve for x.
x = -43/-16
x = 2 11/16

Answer by checkley71(8403)

2(5-4x)+x=3(3x-11) or 10-8x+x=9x-33 or -16x=-43 or x=-43/-16 or x=43/16
proof 2(5-4*43/16)+43/16=3(3*43/16-11) 10-8*43/17+43/17=9*43/16-33 or
(160-344+43)/16=(387-528)/16 or -141/16=-141/16